The coordinates of the points A and B in the x-y plane are (-3,-4) and (5,0) respectively.
(a) Obtain the equation of the circle on AB as diameter, and show that the point (-3,0) lies on the circle.
(b) Find also:
(i) the equation of a line CD in parallel to the line 2x +2y-7=0 whose intercept on the x-axis is -3
(ii) the coordinates of the points of intersection of the line CD with the circle
(iii) the equation of the tangent of the circle at the point (-3,0)
1
Expert's answer
2019-08-25T17:01:30-0400

(a) The center of a circle is the midpoint of a diameter:


"x=\frac{x_A+x_B}{2}=\frac{-3+5}{2}=\frac{2}{2}=1"


"y=\frac{y_A+y_B}{2}=\frac{-4+0}{2}=\frac{-4}{2}=-2"

The coordinates of the centre is (1,-2).

Since the point B lies on the circle, the coordinates (5,0) must satisfy the equation of the circle: "(x-h)^2+(y-k)^2 =r^2", where (h, k) is the center of the circle and r is the radius.


"(5-1)^2 +(0+2)^2 =r^2""16+4=r^2""r^2=20"

The equation of the circle is: "(x-1)^2+(y+2)^2 =20".

Show that the point (-3,0) lies on the circle:


"(-3-1)^2+(0+2)^2 =20""(-4)^2+2^2 =20""16+4=20""20=20"

The coordinates (-3,0) satisfy the equation. The point lies on the circle.


(b) (i) The equation of a line parallel to the given line "2x+2y-7=0" is "2x+2y+k=0", because parallel lines have the same slope and different intercepts.

Since the line CD has x-intercept equal to -3, the point (-3;0) satisfy the equation of CD:


"2\cdot(-3)+2\cdot0+k=0""-6+k=0""k=6"

The equation of the line CD is "2x+2y+6=0".


(ii) Find the coordinates of the points of intersection of the line CD with the circle:


"\begin{cases} 2x+2y+6=0 \\ (x-1)^2 +(y+2)^2 =20 \end{cases}""\begin{cases} y=-x-3 \\ (x-1)^2 +(-x-3+2)^2 =20 \end{cases}""\begin{cases} y=-x-3 \\ (x-1)^2 +(-x-1)^2 =20 \end{cases}""x^2 -2x+1+x^2 +2x+1=20""2x^2 =18""x^2=9""x=3,x=-3"

If "x=3" , then "y=-3-3=-6".

If "x=-3" , then "y=-(-3)-3=3-3=0"

The coordinates of the points of intersection of the line CD with the circle are (3,-6), (-3;0)


(iii) Find  the equation of the tangent of the circle at the point (-3,0). Determine the slope of the radius: "m_r=\frac{y_1-y_2}{x_1-x_2}", where "(x_1,y_1)" is the center of the circle, "(x_2,y_2)" is the point on a circle.



"m_r =\frac{-2-0}{1+3}=\frac{-2}{4}=-\frac{1}{2}"


The slope of the tangent line must be perpendicular to the slope of the radius, so the slope of the line is "m_T=-\frac{1}{m_r}=2"

Use the point/slope form "y-y_1=m_T(x-x_1)" at the point (-3,0):


"y=2(x+3)""y=2x+6"